Find the general solution of the given differential equation

dr / d theta + rsec theta = cos theta
1
Expert's answer
2020-07-20T18:29:08-0400

Question: "\frac {dr}{d\theta}+rsec(\theta)=cos(\theta)"

Solution:

Substituting "r=uv" , "r'=u'v+uv'" : (1)

"u'v+uv'+sec(\theta)uv=cos(\theta)"

"u'v+u(v'+sec(\theta)v)=cos(\theta)" (2)

Setting the part inside () to 0 and separating variables:

"v'+sec(\theta)v=0"

"\frac {dv}{d\theta}=-sec(\theta)v"

"\int \frac {dv}{d\theta}=-\int sec(\theta)"

"ln(v)=-ln(sec(\theta)+tan(\theta)) + ln(k)"

"ln(v)=ln(\frac k {sec(\theta)+tan(\theta)})"

"v=\frac k {sec(\theta)+tan(\theta)}"

Putting "v"  from (3) into equation (2):

"u'\frac k {sec(\theta)+tan(\theta)}=cos(\theta)"

Using trigonometric identities "sec(\theta)=\frac 1 {cos(\theta)}" , "tan(\theta)=\frac {sin(\theta)}{cos(\theta)}" :

"du=\frac 1k cos(\theta)(sec(\theta)+tan(\theta))d\theta=\frac 1k (1+sin(\theta))d\theta"

"u=\frac 1k \int (1+sin(\theta))d\theta = \frac 1k (\theta-cos(\theta)+C)"

Finding r from (1):

"r=uv=\frac {\theta-cos(\theta)+C}{sec(\theta)+tan(\theta)}"

Answer: "r=\frac {\theta-cos(\theta)+C}{sec(\theta)+tan(\theta)}"